I have a question that states: Find the distance between the two plotted points (0,2) and (5,-4). I came up with 5.39 but it is wrong, what would this be?
Answer by checkley75(3666)

What you have, when graphed, is a right triangle with a distance of (5-0)=5 on the x axis & (2+4)=6 on the y axis.
Using the pythagorean theorum we get:
5^2+6^2=c^2
25+36=c^2
61=c^2
c=sqrt61
c=7.81 is the distance between these 2 points.